Use extraction of roots to determine the exact solutions of each quadratic equation:
(2x-1)^2 = 9

Answer by ankor@dixie-net.com(22740)

(2x-1)^2 = 9
:
(2x-1)(2x-1) = 9
FOIL
4x^2 - 4x + 1 = 9
:
4x^2 - 4x + 1 - 9 = 0
:
4x^2 - 4x - 8 = 0
Simplify, divide by 4
4(x^2 - x - 2 ) = 0
Factors to
(x-2)(x-1) = 0
x = +2
x = -1